OpenStudy (anonymous):

what is the slope intercept form of the equation x-3y=3

OpenStudy (tkhunny):

Have you considered solving for y?

OpenStudy (smored):

x-3y=3 -3y=x+3 y=1x/-3-1

OpenStudy (smored):

\[y=\frac{ 1 }{ -3 }x-1\]

OpenStudy (tkhunny):

So close!!! Do that "subtract x" thing again. Something went horribly wrong.

OpenStudy (anonymous):